OpenStudy (anonymous):

find an equation of the line containing (-2,1) with slope 8.

OpenStudy (anonymous):

wrong it's actually y=8x+17

OpenStudy (anonymous):

where is the 17 from?

OpenStudy (anonymous):

y=mx+b is an equation of the line. since we have the slope and coordinate point. we plug in the slope and the x,y point in also. 1=8(-2)+b and solve for b 1=-16+b Since we subtract 1--16=1+16 b=17

OpenStudy (anonymous):

thanks